OpenStudy (anonymous):

How do you put this in point-slope form? through (4,1), parallel to 2x+5y=10

OpenStudy (ribhu):

y-1 = -2/5 (x-4)

OpenStudy (michele_laino):

Hint: please note that the equation of the given line, namely: \[2x + 5y = 10\] can be rewritten as below: \[y = - \frac{2}{5}x + 2\]

OpenStudy (michele_laino):

so, waht is its slope

OpenStudy (anonymous):

2/5 @Michele_Laino

OpenStudy (michele_laino):

no, please it is -2/5

OpenStudy (anonymous):

Yeah I just wrote it how you had it in the equation, it was a positive but my mistake. So now that I have the slope, all I do is plug it into the formula?

OpenStudy (michele_laino):

now, keep in mind that parallel lines have the same slope, so the slope of requested line is m= -2/5, furthermore, you have to apply this equation: \[y - {y_0} = m(x - {x_0})\] where: \[\left( {{x_0},{y_0}} \right) = \left( {4,1} \right)\]

OpenStudy (michele_laino):

so, please substitute those value of x_0, y_0 and m into the above equation

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

I'm kinda confused with how I'm multiplying my fraction. This is how I set my problem up: y-1=-2/5(x-4) distributed: y-1=-2/5x+ ? I know its -2/5 times 4/1. Then I make the denominators alike = -2/5 times 20/5 Am I doing something wrong?

OpenStudy (michele_laino):

your first step is correct your second step is: \[y - 1 = - \frac{2}{5}x + \frac{8}{5}\] since: \[\left( { - \frac{2}{5}} \right) \cdot \left( { - 4} \right) = \frac{8}{5}\]

OpenStudy (anonymous):

Oh I get it now Michele. I was supposed to multiply without like denominators, just straight across. My teacher told me this but I didn't get it

OpenStudy (anonymous):

So it's gonna be \[y-1=-\frac{ 2 }{ 5 }x+\frac{ 8 }{ 5 }\]

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

But how do you subtract 1 from \[\frac{ 8 }{ 5 }\]

OpenStudy (anonymous):

I mean add 1

OpenStudy (michele_laino):

please note that you have to add 1 to both sides of your last equation, namely: \[y - 1 + 1 = - \frac{2}{5}x + \frac{8}{5} + 1\] furthermore, we have this: \[\frac{8}{5} + 1 = \frac{{8 + 1 \cdot 5}}{5} = ...?\]

OpenStudy (anonymous):

\[\frac{ 45 }{ 5 }\]

OpenStudy (michele_laino):

no, please you have to compute the multiplication first

OpenStudy (anonymous):

Oh yeah you right, \[\frac{ 8+(1\times5) }{ 5 } =\frac{ 8+5 }{ 5 } = \frac{ 13 }{ 5 } =2 \frac{ 3 }{ 5 }\]

OpenStudy (michele_laino):

perfect! well done!

OpenStudy (anonymous):

Why did you multiply the top by 5?

OpenStudy (michele_laino):

since the denominator of 1 is 1, so you have to add this fractions: \[\frac{8}{5} + \frac{1}{1}\] now the least common multiple between 5 and 1 is 5, so the denominator of your result is 5: \[\frac{8}{5} + \frac{1}{1} = \frac{?}{5}\] now, since 5:5= 1, then you have to multiply 8 by 1, and since 5:1=5, then you have to multiply 1 by 5, like this: \[\frac{8}{5} + \frac{1}{1} = \frac{{\left( {8 \times 1} \right) + \left( {1 \times 5} \right)}}{5} = \frac{{8 + 5}}{5}\]

OpenStudy (anonymous):

So i'ts kinda like you cross multiply without the equal sign in the middle then. Do you always do this for problems like these?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Okay, thank you Michele, you help me get it :)

OpenStudy (michele_laino):

thank you! :)